Monday 12 7
This presentation is the property of its rightful owner.
Sponsored Links
1 / 8

Monday 12.7 PowerPoint PPT Presentation


  • 105 Views
  • Uploaded on
  • Presentation posted in: General

Monday 12.7. 03:00 Stable beams 9_6_6_6 09:40 back to ADJUST to switch on damper and switch off octupoles . 10:20 back to STABLE BEAMS. 11:57 beam dumped by LHC VELO when switching back to ADJUST mode after ADJUST handshake. VELO was not fully out.

Download Presentation

Monday 12.7

An Image/Link below is provided (as is) to download presentation

Download Policy: Content on the Website is provided to you AS IS for your information and personal use and may not be sold / licensed / shared on other websites without getting consent from its author.While downloading, if for some reason you are not able to download a presentation, the publisher may have deleted the file from their server.


- - - - - - - - - - - - - - - - - - - - - - - - - - E N D - - - - - - - - - - - - - - - - - - - - - - - - - -

Presentation Transcript


Monday 12 7

Monday 12.7

  • 03:00 Stable beams 9_6_6_6

    • 09:40 back to ADJUST to switch on damper and switch off octupoles.

    • 10:20 back to STABLE BEAMS.

    • 11:57 beam dumped by LHC VELO when switching back to ADJUST mode after ADJUST handshake. VELO was not fully out.

    • End of a very good fill, luminosities still at 7E29. No problem with damper in stable beams.

  • 13:40 Access.

    • S12 QPS fixed.

    • FMCM threshold increase on RMSD.B1(2). Up by a factor ~ 4 at injection. Rough test with beam is OK (vertical shift of 3 mm on BTV in front of TDE). To be repeated with a complete OFF to the PC (injection).

    • RF access in P4.

8:30 meeting


Tune spectra damper on off

Tune spectra damper On-Off

8:30 meeting


Monday 12 71

Monday 12.7

  • 10:00-17:00 SPS multi-bunch cycle

    • Continued setting up of the cycle with 4 injections. Emittances below nominal.

    • Adjusted extraction kicker delays to extract ALL 4 bunches.

    • Extraction to upstream TEDs for both B1 and B2.

    • Extraction to TI2 downstream ok.

    • Everything looking ok – next step is injection into the LHC.

  • 15:00 LHCb with - polarity

  • 17:00 TDI P2 problem

    • Does not reach requested setting. Mechanical problem to be fixed today.

    • TDI in position with help of the piquet.

  • 17:30 Injection

    • Large orbit perturbation from LHCb polarity inversion – large local kick at MCBX.1R8.

8:30 meeting


Monday 12 72

Monday 12.7

  • 19:30: Filling 12_8_8_8 9E10/bunch

  • 20:00: Ramp

  • 20:45: Flat top

    • Incorporated RT trims from orbit into settings/functions. Worked partly only – anything changed in the past 2 weeks???

    • When the OFB was switched off, the settings of the RT suddenly jumped to NaN for all vertical orbit correctors (????????), horizontal plane OK.

    • Switching ON/OFF OFB seemed to indicate that it was still working…

  • 21:20 Squeeze

    • Beam lost due to vertical orbit change: the OFB did not act in vertical.

    • Beam dumped by losses in IR7 (84 s RS on MQWA).

  • 23:00 Setup

    • MKI problem IR8 – faulty after pressure rise.

  • 00:30 Injection

8:30 meeting


Tuesday 13 7

Tuesday 13.7

  • 02:00: Again some issue get the OFB ON

  • 02:30: Ramp

  • 04:00: Squeeze

    • OFB off by itself while squeezing. Orbit change moderate and beam survived. Manual correction at 7 m, then OFB on again.

  • 05:20 Stable beams fill 1224

    • L 1.4E30 cm-2s-1.

    • Some beam loss after ATLAS lumi scan.

    • Strong losses correlated with lines at 0.315 in Q spectrum.

8:30 meeting


Fill 1224

Fill 1224

8:30 meeting


Tune spectrum lines

Tune spectrum lines…

8:30 meeting


Tuesday 13 71

Tuesday 13.7

  • Keep this fill for some time, try again damper.

  • Multi-bunch injection studies later today.

  • Fill overnight.

  • Check OFB issues during the day (start with offline checks).

8:30 meeting


  • Login